OpenStudy (anonymous):

the volume of a cube is decreasing at a rate of 0.24 ft^3/min. What is the rate of change of the side length when the side length are 2 feet.

OpenStudy (anonymous):

ok what do we know

OpenStudy (anonymous):

a^3=V

OpenStudy (anonymous):

ko

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

dV/da= 0.24 ft^3/min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which also equals 3a^2

OpenStudy (anonymous):

9a

OpenStudy (anonymous):

? whered you get 9a

OpenStudy (anonymous):

3a^2

OpenStudy (anonymous):

sorry.. i dont follow

OpenStudy (anonymous):

oh wait time has a place in this..sorry

OpenStudy (anonymous):

dV/dt=3a^2(da/dt)

OpenStudy (anonymous):

we have to solve for da/dt at a=2

OpenStudy (anonymous):

do you follow?

OpenStudy (campbell_st):

well it looks like this \[\frac{dV}{dt} = \frac{dV}{dx} \times \frac{dx}{dt}\] you know \[\frac{dV}{dt} = -0.24\] you also know the volume of a cube side length x \[V = x^3\] then \[\frac{dV}{dx} = 3x^2\] so you know by substitution \[-0.24 = 3x^2 \times \frac{dx}{dt}\] substitiute x = 2 and solve for dx/dt

OpenStudy (campbell_st):

the units for the rate of change in the side length is ft/min

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

so then the answer would be -12.24